Let Z have a standard normal N(0,1) distribution.
(a) Given zα defined by P(Z < zα) = α; find z0.879.
(b) Calculate the value of k such that P(−1.61 < Z < k) = 0.8958.
1
Expert's answer
2020-10-12T11:51:46-0400

(a) using table, Z for 0.879 equals 1.17;

(b) P(−1.61 < Z < k) =0.8958.

table(k) - table(−1.61) = 0.8958

table(k) - 0.0537 = 0.8958

table(k) = 0.8958+0.0537=0.9495

from area table

k = 1.64